OpenStudy (anonymous):

Find the direct variation equation which passes through (0, 0) and (4, 1). y = 4x y = -4x y = 1/4 x y = -1/4 x

OpenStudy (anonymous):

You can start by finding its Slope...

OpenStudy (anonymous):

\[Slope = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

1/4 is the slope

OpenStudy (anonymous):

Right.. And equation is given by: y = mx + c where m = Slope and c = y-intercept..

OpenStudy (anonymous):

As this line is passing through origin or the line cuts y-axis at 0, hence c = 0..

OpenStudy (anonymous):

So, your equation reduces to : y = mx You know value of m here, so just put and find your equation.. :)

OpenStudy (anonymous):

So, the answer would be y=1/4x? Since the c is 0?

OpenStudy (anonymous):

Yep.. :)

OpenStudy (anonymous):

Is the y intercept y1?

OpenStudy (anonymous):

What??

OpenStudy (anonymous):

Yes here you can say..

OpenStudy (anonymous):

but how do you find the y intercept? (c)

OpenStudy (anonymous):

Basically, y-intercept is the coordinate or length on y-axis where your line cuts y-axis.

OpenStudy (anonymous):

Your given points are : (0,0) and (4,1) so try to graph them: 

OpenStudy (anonymous):

You can find it mathematically also, here is the method: Suppose you are given with points : (3,4) and (4,2) Find Slope first, Slope(m) = (2-4)/(4-3) = -2

OpenStudy (anonymous):

Now, you know : y = mx + c And m = -2: so just put here: so you get: y = -2x + c Now, you should know this, if you know a point lying on your line, then that point should satisfy the equation of Line, I mean if line is passing through (2,4) then this point(2.4) should satisfy the equation of that line..

OpenStudy (anonymous):

So, (3,4) point must satisfy your line equation.. y = -2x+c Substitute that point here in this equation: For point (3,4) : 4 = -2(3) + c \(c = 10\)

OpenStudy (anonymous):

Similarly if you will put second point (4,2) in your equation, then also you must get the same value of c, try it..

OpenStudy (anonymous):

Getting??

OpenStudy (anonymous):

Dang, now I get it. This helped so much!!! Thank you tons

OpenStudy (anonymous):

It is okay.. :)